# On the Robustness of Kernel Goodness-of-Fit Tests

### v3

@@ Appendix B Complementary Experiments
This section includes additional experimental results. Section B.1 discusses the rate requirement in Theorem 3 and demonstrates that the KSD test is no longer qualitatively robust when this condition is not met. Section B.2 reviews the robust MMD tests included in Section 5.1 and presents implementation details. Section B.4 includes an ablation study on the choice of kernel bandwidths in the robust-KSD test. Section B.6 studies its scalability with dimension. Section B.8 compares robust-KSD using two different bootstrap methods: the weighted bootstrap used throughout this work, and a wild bootstrap that is more prevalent in the kernel testing literature.
@@ B.1 Decay Rate of Contamination Ratio
We show empirically that the rate requirement in the qualitative robustness result Theorem 3 is not an artifact of the proof, i.e., the tilted-KSD test is no longer qualitatively robust to Huber’s contamination models with contamination ratio $\epsilon_{n}$ if $\epsilon_{n}=n^{-r}$ for any $r\leq 1/2$, where $n$ is the sample size. We run the standard KSD test with a tilted kernel under the contaminated Gaussian model as described in Section 5.1 with dimension $d=1$ and outlier $z=10$. The contamination ratio is chosen to be $\epsilon_{n}=n^{-r}$ for different choices of $r$, and the probability of rejection as $n$ grows is shown in Figure 9. The results are averaged over $400$ repetitions instead of 100 repetitions as in Section 5.1 to reduce numerical uncertainty. When $r>0.5$, the rejection probability converges to the prescribed test level $\alpha$, which is the limit of the rejection probability without contamination. This aligns with Theorem 3. However, this rejection probability no longer converges to $\alpha$ when $r\leq 0.5$, thus showing that the tilted-KSD test is no longer qualitatively robust.
@@ B.2 Implementation Details for the Robust MMD Tests
This section provides implementation details for the two robust MMD-based tests included in Section 5.1. Both tests use the Maximum Mean Discrepancy (MMD; Gretton et al.,, 2012, Definition 2) between the data-generating distribution $Q$ and the model $P$ as a measure of their disparity. Given a reproducing kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$, the squared MMD between $Q$ and $P$ are defined as
@@
Given independent samples $\mathbb{X}_{n}=\{\mathbf{X}_{i}\}_{i=1}^{n}$ from $Q$ and $\mathbb{Y}_{m}=\{\mathbf{Y}_{j}\}_{j=1}^{m}$ from $P$, the MMD can be estimated directly by $D_{\mathrm{MMD}}^{2}(Q_{n},P_{m})$, where $Q_{n}$ and $P_{m}$ are the empirical measures formed by $\mathbb{X}_{n}$ and $\mathbb{Y}_{m}$, respectively (Gretton et al.,, 2012). The computational cost of such estimate scales with sample size $n$ as $\mathcal{O}((n+m)^{2}+mC_{\text{sim}})$, where $C_{\text{sim}}$ is the cost of simulating one datum from $P$, compared with $\mathcal{O}(n^{2}+nC_{\text{score}})$ for KSD, where $C_{\text{score}}$ denotes the cost of a single score evaluation. In our experiment in Section 5.1, $P$ is a Gaussian model, so both simulation and score evaluation are trivial and $C_{\text{sim}},C_{\text{score}}$ are both small. Hence, we set $m=n$ so that the costs of MMD and of KSD are comparable. Moreover, we will use the squared-exponential kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\exp(-\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}/(2\gamma^{2}))$ and set $\gamma^{2}$ using the median heuristic. Squared-exponential kernels are popular for MMDs and are also used in the original papers that proposed the two robust MMD tests (Sun and Zou,, 2023; Schrab and Kim,, 2024).
@@ B.2.1 The MMD-Dev Test
Proof of Lemma 23  For any $Q\in\mathcal{P}(\mathbb{R}^{d})$, we define its kernel mean embedding (Berlinet and Thomas-Agnan,, 2004, Chapter 4) as $\mu_{Q}(\cdot)\coloneqq\mathbb{E}_{\mathbf{X}\sim Q}[k(\mathbf{X},\cdot)]$, which is well-defined since $k$ is bounded. By Sriperumbudur et al., (2010, Theorem 1), the MMD between any $Q,P\in\mathcal{P}(\mathbb{R}^{d})$ can be equivalently written as $D_{\mathrm{MMD}}(Q,P)=\|\mu_{Q}-\mu_{P}\|_{\mathcal{H}_{k}}$, where $\mathcal{H}_{k}$ is the RKHS associated with $k$. Pick $\epsilon_{0}\in[0,1]$ and $R\in\mathcal{P}$. For any $\epsilon\in[0,\epsilon_{0}]$,
@@ B.2.2 The dcMMD Test
$H_{0}^{\mathrm{dcMMD}}:\textrm{ At least $(1-\epsilon_{0})\times n$ random variables in $\mathbb{X}_{n}$ are i.i.d.\ from $P$.}$
@@
Figure 10: Left. Estimated $\tau_{\infty}$ using the trick described in Section 4.2 compared with the ground-truth; the estimate becomes more accurate with larger samples. Right. Rejection probability of robust-KSD under an outlier-contaminated Gaussian model; the Type-I error remains calibrated even for small sample sizes.
@@ B.3 Approximating the Supremum of the Stein Kernel
As described in Section 4.2, in our experiments we approximated the intractable supremum $\tau_{\infty}=\sup_{\mathbf{x}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x})$ by $\max_{i=1,\ldots,n}u_{p}(\mathbf{x}_{i},\mathbf{x}_{i})$. We now study how this approximation affects the performance of the robust-KSD test. We run the same contaminated Gaussian experiment in Section 5.1 with $d=1$ and outlier location $z=10$. For this model, $\mathbf{x}\mapsto u_{p}(\mathbf{x},\mathbf{x})$ is a simple, univariate function, so we can compute the ground-truth $\tau_{\infty}$ accurately by numerical optimization. Clearly, the ground truth $\tau_{\infty}$ is equivalent to taking $n=\infty$. This is evidenced by the left plot in Figure 10, which shows that the finite-sample approximation $\max_{i=1,\ldots,n}u_{p}(\mathbf{x}_{i},\mathbf{x}_{i})$ becomes more accurate as $n$ increases. The right plot of Figure 10 shows the rejection probability of robust-KSD with $\theta$ set to control at most $\epsilon_{0}=0.05$ proportion of contamination. Remarkably, even when this approximation under-estimates $\tau_{\infty}$ (which is the case for small $n$), robust-KSD still remains well-calibrated. This is not surprising: the supremum $\tau_{\infty}$ represents the maximal contribution of a single datum $\mathbf{x}$ taking arbitrary values in the entire sample space $\mathbb{R}^{d}$, regardless of whether it is present in the observed data set. However, in practice, it suffices to control the contribution of any single datum in the observed data set.
@@ B.4 Ablation Study for Kernel Bandwidth in the Standard KSD Test
We evaluate how the choice of the kernel bandwidth affects the performance of the standard KSD test. We run the same contaminated Gaussian experiment as in the previous subsection. The standard KSD test uses an IMQ kernel with bandwidth $\lambda^{2}\in\Lambda\cup\{\lambda_{\mathrm{med}}^{2}\}$, where $\Lambda=\{0.01,0.1,1,10,100\}$ and $\lambda_{\mathrm{med}}^{2}$ is the bandwidth chosen by median heuristic. We also include KSDAgg, which also uses an IMQ kernel for a fair comparison. As shown in Figure 9, all standard IMQ-KSD tests reject the point null with high probability for large outlier values, regardless of the bandwidth value. This suggests that no fixed bandwidth can ensure robustness, which is consistent with Theorem 1, and is because the Stein kernel is unbounded regardless of the choice of $\lambda$, thus the non-robustness issue persists. The KSDAgg test is even more sensitive to contamination than the IMQ-KSD tests. As noted in Section 5.1, this is because KSDAgg is designed to optimally combine multiple bandwidths to boost the test power against all alternatives, including contaminated models.
@@ B.5 Gaussian Mean-Shift Experiment
We run a Gaussian mean-shift experiment to demonstrate the performance of our robust test against model deviations other than contamination. We use a Gaussian model $\mathcal{N}(0,I_{d})$ in dimension $d=50$, and draw data from $Q_{\mu_{0}}=\mathcal{N}(\mu_{0}e_{1},I_{d})$, where $e_{1}=(1,0,\ldots,0)^{\top}\in\mathbb{R}^{d}$ and $\mu_{0}\in\mathbb{R}$ is a mean-shift. The uncertainty radius $\theta$ is chosen to be the Tilted-KSD value corresponding to $\mu_{0}=0,0.2$ and $0.6$, respectively. The purpose of this experiment is to demonstrate that the robust-KSD test is well-calibrated when $D(Q_{\mu_{0}},P)\leq\theta$ and consistent when $D(Q_{\mu_{0}},P)>\theta$. As shown in Figure 11, the standard tests reject with probability higher than robust-KSD. This is again expected since the standard tests are not robust to contamination. For mean-shift values not greater than the black vertical line, $Q_{\mu_{0}}\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$, so we are under the null hypothesis and robust-KSD rejects no more frequently than the level, showing that it is well-calibrated. For larger $\mu_{0}$, robust-KSD rejects with probability approaching one, thus showing its power. Moreover, when $\mu_{0}=0$ so that $\theta=0$, robust-KSD becomes identical to the standard Tilted-KSD test, showing that robust-KSD is indeed a generalization of the standard test.
@@ B.6 Scalability with Dimension
We run the Gaussian mean-shift experiment in Section B.5 in different dimensions; other experimental setups remain unchanged. The uncertainty radius for R-KSD is chosen to be the (non-squared) KSD value corresponding to mean-shift $\mu=0.3$. Results are reported in Figure 12. As dimension increases, both the two standard tests and R-KSD have declining power in correctly rejecting for large values of $\mu$. This shows that both the standard and the robust KSD tests suffer from the curse-of-dimensionality, a known issue for kernel-based tests (Huang et al.,, 2023; Reddi et al.,, 2015; Ramdas et al.,, 2015). Unsurprisingly, this issue is more prominent for the R-KSD test. This is because the KSD-ball $\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ could potentially include more distributions in higher dimensions.
@@ B.7 Different Contamination Distributions
In most of our experiments in Section 5, the data-generating distribution takes the form $Q=(1-\epsilon_{0})P+\epsilon_{0}R$ with the contamination distribution $R=\delta_{\mathbf{z}}$ being a Dirac delta taking a single value at $\mathbf{z}$. We now investigate how the choice of $R$ affects the empirical performance of R-KSD. Importantly, our results in Section 4 make no assumption on the form of $R$.
We set $P=\mathcal{N}(0,1)$ and generate data from $Q=(1-\epsilon)P+\epsilon\mathcal{N}(10,\sigma^{2})$ with varying $\sigma>0$. Following Proposition 5, we choose $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$, and fix the contamination tolerance to be $\epsilon_{0}=0.05$. Since Proposition 5 holds for all $R$, we expect R-KSD to remain (asymptotically) valid in this setting.
Results are shown in Figure 13. For larger $\sigma$, all tests, including our R-KSD, saw a exhibit reduced test power. This is because, when $\sigma$ is large, the noise component $R$ becomes more dispersed, making it harder for KSD to detect discrepancies. This can be confirmed by the rightmost plot, which shows that the KSD between $Q$ and $P$ decreases with $\sigma$. Crucially, our R-KSD test is able to control the Type-I error regardless of the value of $\sigma$.
Figure 13: Mixture-of-Gaussian experiment. Left. Probability of rejection under different standard deviations for the contamination component. Right. KSD estimates. Grey dotted line. Test level $\alpha=0.05$. Black dash-dot line. Uncertainty radius $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$, with $\tau_{\infty}$ estimated following Section 4.2.
@@ B.8 Weighted Bootstrap and Wild Bootstrap
We compare the robust-KSD test using the weighted bootstrap described in Section 2.2 (equivalent to the Efron’s bootstrap) against the wild bootstrap due to Leucht and Neumann, (2013). For the standard KSD test, weighted bootstrap was used in Liu et al., (2016), while the wild bootstrap is more popular in the literature (Chwialkowski et al.,, 2016; Schrab et al.,, 2022; Liu et al.,, 2023). Compared with weighted bootstrap, the wild approach is more flexible as it can work for dependent samples (Chwialkowski et al.,, 2014), and theoretical guarantees on the power of the standard test were only proved with the wild bootstrap (Schrab et al.,, 2022). In this work, we have used the weighted approach since we can then leverage existing theoretical results to show the validity of the test; see Section A.5.1. This is for simplicity rather than necessity.
We numerically compare these two bootstrap methods using the contaminated Gaussian model in Section 5.1 and show that the use of weighted bootstrap does not negatively impact the test. Figure 14 shows the rejection probability of the robust-KSD test using the weighted bootstrap (R-KSD) and wild bootstrap (R-KSD-Wild). The uncertainty radius is chosen to be $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$ for different values of $\epsilon_{0}$, following Section 5.1. In particular, $\epsilon_{0}=0$ corresponds to setting $\theta=0$, in which case robust-KSD reduces to the standard KSD. As evident from the plot, these two bootstrap methods produce almost identical results regardless of the value of $\theta$.
Figure 14: Probability of rejection under an outlier-contaminated Gaussian model using the weighted bootstrap and the wild bootstrap. Black dash-dot vertical line. Uncertainty radius $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$, with $\tau_{\infty}$ estimated following Section 4.2.

### v4

@@ Appendix B Complementary Experiments
This section includes additional experimental results. Section˜B.1 discusses the rate requirement in Theorem˜3 and demonstrates that the KSD test is no longer qualitatively robust when this condition is not met. Section˜B.2 reviews the robust MMD tests included in Section˜5.1 and presents implementation details. Section˜B.4 includes an ablation study on the choice of kernel bandwidths in the robust-KSD test. Section˜B.6 studies its scalability with dimension. Section˜B.8 compares robust-KSD using two different bootstrap methods: the weighted bootstrap used throughout this work, and a wild bootstrap that is more prevalent in the kernel testing literature.
@@ B.1 Decay Rate of Contamination Ratio
We show empirically that the rate requirement in the qualitative robustness result Theorem˜3 is not an artifact of the proof, i.e., the tilted-KSD test is no longer qualitatively robust to Huber’s contamination models with contamination ratio $\epsilon_{n}$ if $\epsilon_{n}=n^{-r}$ for any $r\leq 1/2$, where $n$ is the sample size. We run the standard KSD test with a tilted kernel under the contaminated Gaussian model as described in Section˜5.1 with dimension $d=1$ and outlier $z=10$. The contamination ratio is chosen to be $\epsilon_{n}=n^{-r}$ for different choices of $r$, and the probability of rejection as $n$ grows is shown in Figure˜9. The results are averaged over $400$ repetitions instead of 100 repetitions as in Section˜5.1 to reduce numerical uncertainty. When $r>0.5$, the rejection probability converges to the prescribed test level $\alpha$, which is the limit of the rejection probability without contamination. This aligns with Theorem˜3. However, this rejection probability no longer converges to $\alpha$ when $r\leq 0.5$, thus showing that the tilted-KSD test is no longer qualitatively robust.
@@ B.2 Implementation Details for the Robust MMD Tests
This section provides implementation details for the two robust MMD-based tests included in Section˜5.1. Both tests use the Maximum Mean Discrepancy (MMD; Gretton et al.,, 2012, Definition 2) between the data-generating distribution $Q$ and the model $P$ as a measure of their disparity. Given a reproducing kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$, the squared MMD between $Q$ and $P$ are defined as
@@
Given independent samples $\mathbb{X}_{n}=\{\mathbf{X}_{i}\}_{i=1}^{n}$ from $Q$ and $\mathbb{Y}_{m}=\{\mathbf{Y}_{j}\}_{j=1}^{m}$ from $P$, the MMD can be estimated directly by $D_{\mathrm{MMD}}^{2}(Q_{n},P_{m})$, where $Q_{n}$ and $P_{m}$ are the empirical measures formed by $\mathbb{X}_{n}$ and $\mathbb{Y}_{m}$, respectively (Gretton et al.,, 2012). The computational cost of such estimate scales with sample size $n$ as $\mathcal{O}((n+m)^{2}+mC_{\text{sim}})$, where $C_{\text{sim}}$ is the cost of simulating one datum from $P$, compared with $\mathcal{O}(n^{2}+nC_{\text{score}})$ for KSD, where $C_{\text{score}}$ denotes the cost of a single score evaluation. In our experiment in Section˜5.1, $P$ is a Gaussian model, so both simulation and score evaluation are trivial and $C_{\text{sim}},C_{\text{score}}$ are both small. Hence, we set $m=n$ so that the costs of MMD and of KSD are comparable. Moreover, we will use the squared-exponential kernel $k(\mathbf{x},\mathbf{x}^{\prime})=\exp(-\|\mathbf{x}-\mathbf{x}^{\prime}\|_{2}^{2}/(2\gamma^{2}))$ and set $\gamma^{2}$ using the median heuristic. Squared-exponential kernels are popular for MMDs and are also used in the original papers that proposed the two robust MMD tests (Sun and Zou,, 2023; Schrab and Kim,, 2024).
@@ B.2.1 The MMD-Dev Test
Proof of Lemma˜23  For any $Q\in\mathcal{P}(\mathbb{R}^{d})$, we define its kernel mean embedding (Berlinet and Thomas-Agnan,, 2004, Chapter 4) as $\mu_{Q}(\cdot)\coloneqq\mathbb{E}_{\mathbf{X}\sim Q}[k(\mathbf{X},\cdot)]$, which is well-defined since $k$ is bounded. By Sriperumbudur et al., (2010, Theorem 1), the MMD between any $Q,P\in\mathcal{P}(\mathbb{R}^{d})$ can be equivalently written as $D_{\mathrm{MMD}}(Q,P)=\|\mu_{Q}-\mu_{P}\|_{\mathcal{H}_{k}}$, where $\mathcal{H}_{k}$ is the RKHS associated with $k$. Pick $\epsilon_{0}\in[0,1]$ and $R\in\mathcal{P}$. For any $\epsilon\in[0,\epsilon_{0}]$,
@@ B.2.2 The dcMMD Test
[…]
@@
Figure 10: Left. Estimated $\tau_{\infty}$ using the trick described in Section˜4.2 compared with the ground-truth; the estimate becomes more accurate with larger samples. Right. Rejection probability of robust-KSD under an outlier-contaminated Gaussian model; the Type-I error remains calibrated even for small sample sizes.
@@ B.3 Approximating the Supremum of the Stein Kernel
As described in Section˜4.2, in our experiments we approximated the intractable supremum $\tau_{\infty}=\sup_{\mathbf{x}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x})$ by $\max_{i=1,\ldots,n}u_{p}(\mathbf{x}_{i},\mathbf{x}_{i})$. We now study how this approximation affects the performance of the robust-KSD test. We run the same contaminated Gaussian experiment in Section˜5.1 with $d=1$ and outlier location $z=10$. For this model, $\mathbf{x}\mapsto u_{p}(\mathbf{x},\mathbf{x})$ is a simple, univariate function, so we can compute the ground-truth $\tau_{\infty}$ accurately by numerical optimization. Clearly, the ground truth $\tau_{\infty}$ is equivalent to taking $n=\infty$. This is evidenced by the left plot in Figure˜10, which shows that the finite-sample approximation $\max_{i=1,\ldots,n}u_{p}(\mathbf{x}_{i},\mathbf{x}_{i})$ becomes more accurate as $n$ increases. The right plot of Figure˜10 shows the rejection probability of robust-KSD with $\theta$ set to control at most $\epsilon_{0}=0.05$ proportion of contamination. Remarkably, even when this approximation under-estimates $\tau_{\infty}$ (which is the case for small $n$), robust-KSD still remains well-calibrated. This is not surprising: the supremum $\tau_{\infty}$ represents the maximal contribution of a single datum $\mathbf{x}$ taking arbitrary values in the entire sample space $\mathbb{R}^{d}$, regardless of whether it is present in the observed data set. However, in practice, it suffices to control the contribution of any single datum in the observed data set.
@@ B.4 Ablation Study for Kernel Bandwidth in the Standard KSD Test
We evaluate how the choice of the kernel bandwidth affects the performance of the standard KSD test. We run the same contaminated Gaussian experiment as in the previous subsection. The standard KSD test uses an IMQ kernel with bandwidth $\lambda^{2}\in\Lambda\cup\{\lambda_{\mathrm{med}}^{2}\}$, where $\Lambda=\{0.01,0.1,1,10,100\}$ and $\lambda_{\mathrm{med}}^{2}$ is the bandwidth chosen by median heuristic. We also include KSDAgg, which also uses an IMQ kernel for a fair comparison. As shown in Figure˜9, all standard IMQ-KSD tests reject the point null with high probability for large outlier values, regardless of the bandwidth value. This suggests that no fixed bandwidth can ensure robustness, which is consistent with Theorem˜1, and is because the Stein kernel is unbounded regardless of the choice of $\lambda$, thus the non-robustness issue persists. The KSDAgg test is even more sensitive to contamination than the IMQ-KSD tests. As noted in Section˜5.1, this is because KSDAgg is designed to optimally combine multiple bandwidths to boost the test power against all alternatives, including contaminated models.
@@ B.5 Gaussian Mean-Shift Experiment
We run a Gaussian mean-shift experiment to demonstrate the performance of our robust test against model deviations other than contamination. We use a Gaussian model $\mathcal{N}(0,I_{d})$ in dimension $d=50$, and draw data from $Q_{\mu_{0}}=\mathcal{N}(\mu_{0}e_{1},I_{d})$, where $e_{1}=(1,0,\ldots,0)^{\top}\in\mathbb{R}^{d}$ and $\mu_{0}\in\mathbb{R}$ is a mean-shift. The uncertainty radius $\theta$ is chosen to be the Tilted-KSD value corresponding to $\mu_{0}=0,0.2$ and $0.6$, respectively. The purpose of this experiment is to demonstrate that the robust-KSD test is well-calibrated when $D(Q_{\mu_{0}},P)\leq\theta$ and consistent when $D(Q_{\mu_{0}},P)>\theta$. As shown in Figure˜11, the standard tests reject with probability higher than robust-KSD. This is again expected since the standard tests are not robust to contamination. For mean-shift values not greater than the black vertical line, $Q_{\mu_{0}}\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$, so we are under the null hypothesis and robust-KSD rejects no more frequently than the level, showing that it is well-calibrated. For larger $\mu_{0}$, robust-KSD rejects with probability approaching one, thus showing its power. Moreover, when $\mu_{0}=0$ so that $\theta=0$, robust-KSD becomes identical to the standard Tilted-KSD test, showing that robust-KSD is indeed a generalization of the standard test.
@@ B.6 Scalability with Dimension
We run the Gaussian mean-shift experiment in Section˜B.5 in different dimensions; other experimental setups remain unchanged. The uncertainty radius for R-KSD is chosen to be the (non-squared) KSD value corresponding to mean-shift $\mu=0.3$. Results are reported in Figure˜12. As dimension increases, both the two standard tests and R-KSD have declining power in correctly rejecting for large values of $\mu$. This shows that both the standard and the robust KSD tests suffer from the curse-of-dimensionality, a known issue for kernel-based tests (Huang et al.,, 2023; Reddi et al.,, 2015; Ramdas et al.,, 2015). Unsurprisingly, this issue is more prominent for the R-KSD test. This is because the KSD-ball $\mathcal{B}^{\mathrm{KSD}}(P;\theta)$ could potentially include more distributions in higher dimensions.
@@ B.7 Different Contamination Distributions
In most of our experiments in Section˜5, the data-generating distribution takes the form $Q=(1-\epsilon_{0})P+\epsilon_{0}R$ with the contamination distribution $R=\delta_{\mathbf{z}}$ being a Dirac delta taking a single value at $\mathbf{z}$. We now investigate how the choice of $R$ affects the empirical performance of R-KSD. Importantly, our results in Section˜4 make no assumption on the form of $R$.
We set $P=\mathcal{N}(0,1)$ and generate data from $Q=(1-\epsilon)P+\epsilon\mathcal{N}(10,\sigma^{2})$ with varying $\sigma>0$. Following Proposition˜5, we choose $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$, and fix the contamination tolerance to be $\epsilon_{0}=0.05$. Since Proposition˜5 holds for all $R$, we expect R-KSD to remain (asymptotically) valid in this setting.
Results are shown in Figure˜13. For larger $\sigma$, all tests, including our R-KSD, saw a exhibit reduced test power. This is because, when $\sigma$ is large, the noise component $R$ becomes more dispersed, making it harder for KSD to detect discrepancies. This can be confirmed by the rightmost plot, which shows that the KSD between $Q$ and $P$ decreases with $\sigma$. Crucially, our R-KSD test is able to control the Type-I error regardless of the value of $\sigma$.
Figure 13: Mixture-of-Gaussian experiment. Left. Probability of rejection under different standard deviations for the contamination component. Right. KSD estimates. Grey dotted line. Test level $\alpha=0.05$. Black dash-dot line. Uncertainty radius $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$, with $\tau_{\infty}$ estimated following Section˜4.2.
@@ B.8 Weighted Bootstrap and Wild Bootstrap
We compare the robust-KSD test using the weighted bootstrap described in Section˜2.2 (equivalent to the Efron’s bootstrap) against the wild bootstrap due to Leucht and Neumann, (2013). For the standard KSD test, weighted bootstrap was used in Liu et al., (2016), while the wild bootstrap is more popular in the literature (Chwialkowski et al.,, 2016; Schrab et al.,, 2022; Liu et al.,, 2023). Compared with weighted bootstrap, the wild approach is more flexible as it can work for dependent samples (Chwialkowski et al.,, 2014), and theoretical guarantees on the power of the standard test were only proved with the wild bootstrap (Schrab et al.,, 2022). In this work, we have used the weighted approach since we can then leverage existing theoretical results to show the validity of the test; see Section˜A.5.1. This is for simplicity rather than necessity.
We numerically compare these two bootstrap methods using the contaminated Gaussian model in Section˜5.1 and show that the use of weighted bootstrap does not negatively impact the test. Figure˜14 shows the rejection probability of the robust-KSD test using the weighted bootstrap (R-KSD) and wild bootstrap (R-KSD-Wild). The uncertainty radius is chosen to be $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$ for different values of $\epsilon_{0}$, following Section˜5.1. In particular, $\epsilon_{0}=0$ corresponds to setting $\theta=0$, in which case robust-KSD reduces to the standard KSD. As evident from the plot, these two bootstrap methods produce almost identical results regardless of the value of $\theta$.
Figure 14: Probability of rejection under an outlier-contaminated Gaussian model using the weighted bootstrap and the wild bootstrap. Black dash-dot vertical line. Uncertainty radius $\theta=\epsilon_{0}\tau_{\infty}^{1/2}$, with $\tau_{\infty}$ estimated following Section˜4.2.